​Café Michigan's​ manager, Gary​ Stark, suspects that demand for mocha latte coffees depends on the price being charged. Based on historical​ observations, Gary has gathered the following​ data, which show the numbers of these coffees sold over six different price​ values:                                                                             Price Number Sold ​$2.70 760 ​$3.50 510 ​$2.00 980 ​$4.20 250 ​$3.10 320 ​$4.05 480 Using simple linear regressionLOADING... and given that the price per cup is ​$2.80​, the forecasted demand for mocha latte coffees will be nothing cups ​(enter your response rounded to one decimal​ place).
Answer #1

Using excel for performing regression with number sold as the dependent variable and price as independent variables.

In excel ,go to data tab-->data analysis and select regression

In the regression dialog select the number sold values as Y range and price values as X range and click OK.

We get the regression output with coefficients of price and intercept .

The trend equation is number sold =Intercept+Coefficient of price *price

Linear trend line is , number sold=1454.60-277.62*price

When price =$ 2.80

Number sold =1454.60-277.62*price=1454.60-277.62*2.80

=677.2

Forecasted demand for latte coffees=677.2 cups

Calculations are as shown below :